How many strings of length 4 on the alphabet {A,B,C,D} do not contain AB as a substring ?
1
Expert's answer
2021-10-11T15:10:55-0400

How many strings of length 4 can be generated from the letters {A,B,C,D} ?


"4^4=256"

If the element "AB" is on the first place


"4^2=16"

If the element "AB" is on the second place


"4^2=16"

If the element "AB" is on the third place


"4^2=16"

How many strings of length 4 on the alphabet {A,B,C,D} do not contain AB as a substring ?


"256-(16+16+16)=208"